# Rigidity of the Borell-Brascamp-Lieb Inequality on Weighted Riemannian Manifolds

### 01 Apr 2026 v1

@@ Abstract
In this paper we discuss some results regarding the rigidity of the Borell-Brascamp-Lieb inequality and the Brunn-Minkowski inequality. We show a theorem of rigidity on curvature and measure of the Borell-Brascamp-Lieb inequality, a generalisation of the curvature rigidity theorem by Balogh-Kristály in [2] to the weighted setting. We present some rigidity results of the Brunn-Minkowski inequality and a few further open problems.
@@ 1 Introduction
Since the seminal works on the application of optimal transport in geometric analysis including [19] by Otto-Villani, [20] by Sturm and [13] by Lott-Villani, it became possible to study a wide range of analytical problems on Riemannian manifolds using the tool of optimal transport. In particular, the theory of optimal transport has been a powerful tool in the study of functional and geometrical inequalities on Riemannian manifolds with curvature and dimension bounds.
First proved in 1952 by Henstock and MacBeth on $\mathbb{R}$ in [12] and later developed by Borell and Brascamp-Lieb in [3] and [4], the Borell-Brascamp-Lieb inequality is an integral inequality concerned with the $L^{1}$-norm of two functions and that of the mean of them.
@@
The Borell-Brascamp-Lieb inequality was proved on Riemannian manifolds by Cordero-Erausquin, McCann and Schmuckenschläger in [6]. It was then generalised further to metric measure spaces by Bacher in [1]. The Borell-Brascamp-Lieb inequality implies a number of functional and geometric inequalities including the Brunn-Minkowski inequality and the Prékopa-Leindler inequality, a diagram showing their relationship can be seen in [11].
Rigidity of the Borell-Brascamp-Lieb inequality on $\mathbb{R}^{n}$ has been studied since 1970s, various papers including [8] by Dubuc since then have discussed this problem. Rigidity of it on (unweighted) Riemannian manifolds with curvature bounded from below was shown by Balogh-Kristály in [2].
This paper generalises the curvature rigidity result of the Borell-Brascamp-Lieb inequality on Riemannian manifolds by Balogh and Kristály, Theorem 4.1 of [2], to weighted Riemannian manifolds. The main result of this paper states that the rigidity of the Borell-Brascamp-Lieb inequality implies identical sectional curvatures and certain behaviour of the measure.
@@
On a weighted Riemannian manifold $(M,g,m=e^{-\psi}\mathrm{vol}_{g})$, satisfying $\mathrm{Ric}_{m,N}\geq Kg$ where $n\leq N<\infty$ and $K\in\mathbb{R}$, equality of the Borell-Brascamp-Lieb inequality implies that the $N$-Ricci curvature along the geodesic of transport $\gamma(t):=\mathcal{F}_{t}(x)$ equals $K$ and the sectional curvatures along the geodesic $\gamma$ are identically $\frac{K}{N-1}$. The measure behaves as an (Nn)(N-n)-th power of a linear combination of trignometric functions along the geodesic $\gamma(t)$ when $K>0$; as an (Nn)(N-n)-degree polynomial of $t$ along $\gamma(t)$ when $K=0$; and as an (Nn)(N-n)-th power of of a linear combination of hyperbolic functions along $\gamma(t)$ when $K<0$ for some constants $c>0$.
@@ 2 Preliminaries
Throughout the paper, we consider a smooth weighted Riemannian manifold $(M,g,m)$ of dimension $n$, where $m=e^{-\psi}\mathrm{vol}_{g}$, satisfying the condition $\mathrm{Ric}_{m,N}(\cdot)\geq Kg(\cdot,\cdot)$, $N\in[n,\infty)$ and $K\in\mathbb{R}$. Recall that the $N$-Ricci curvature $\mathrm{Ric}_{m,N}$ is defined to be:
@@
By the theorem of Brenier-McCann, given two measures on $(M,g,m)$, $\mu_{0}=\rho_{0}m=\rho_{0}e^{-\psi}\mathrm{vol}_{g}$ and $\mu_{1}=\rho_{1}m=\rho_{1}e^{-\psi}\mathrm{vol}_{g}$ in $\mathcal{P}^{2}(M)$ that are absolutely continuous with respect to $m$, there is a unique optimal transference plan from $\mu_{0}$ to $\mu_{1}$, given by $\mathcal{F}_{t}$, $\mu_{t}=(\mathcal{F}_{t})_{\#}\mu_{0}$, the push-forward of $\mu_{0}$ by $\mathcal{F}_{t}(x):=\exp_{x}(t\nabla f(x))$, and $f:M\rightarrow\mathbb{R}$ is a $\frac{d^{2}}{2}$-convex function. We refer to [21] for the fundamentals of optimal transport theory, see also [15] and [10] for the theory of optimal transport on Riemannian manifolds.
We consider the weighted Jacobian $J^{\psi}_{t}(x):=e^{\psi(x)-\psi(\mathcal{F}_{t}(x))}\|(D\mathcal{F}_{t}(x))\|$ along the geodesic $\gamma(t):=\mathcal{F}_{t}(x)$. We have the following concavity condition on the weighted Jacobian:
@@
The interpolation inequality (3) on unweighted Riemannian manifolds was shown in the proof of Theorem 1.7 of [20], and its Finsler version was shown in the proof of Theorem 1.2 of [18]. Here we present a proof on weighted Riemannian manifolds satisfying $\mathrm{Ric}_{m,N}\geq Kg$, along the computation shown in the proof of Theorem 1.7 of [20], with an addition of the factor due to measure to the proof in [20], for later use. See also pages 393-401 of [22].
@@
On a weighted Riemannian manifold $(M,g,m=e^{-\psi(x)}\mathrm{vol}_{g})$ satisfying $\mathrm{Ric}_{m,N}\geq Kg$, the weighted Jacobian $J^{\psi}_{t}(x):=e^{\psi(x)-\psi(\mathcal{F}_{t}(x))}\|(D\mathcal{F}_{t}(x))\|$ defined along the geodesic $\gamma(t):=\mathcal{F}_{t}(x)$, where $\mathcal{F}_{t}(x)=\exp_{x}(t\nabla f(x))$ described as above satisfies the concavity condition
$J^{\psi}_{t}(x)^{\frac{1}{N}}\geq(1-t)\beta^{1-t}_{K,N}(d(x,\mathcal{F}_{1}(x)))^{\frac{1}{N}}+t\beta^{t}_{K,N}(d(x,\mathcal{F}_{1}(x)))^{\frac{1}{N}}J^{\psi}_{1}(x)^{\frac{1}{N}}.$
@@
When $\gamma(t)$ is nonconstant, at $x$, we consider the tangent space $T_{x}M=T_{\mathcal{F}_{0}(x)}M$ and an orthonormal frame $\{e_{0}^{i}\}_{i=1}^{n}$, where we denote $e^{1}_{0}$ to be the unit tangent of the geodesic: $\frac{\gamma^{\prime}(0)}{\|\gamma^{\prime}(0)\|}$. We can thus define n1n-1 Jacobi fields $J_{i}(t)=\frac{\partial}{\partial s}\exp_{x}(t(\nabla f(x)+se^{i}_{0}))$ for $i\in\{2,...,n\}$ that are orthogonal to $\gamma^{\prime}(t)$, and the vectors $\{J_{i}(t)\}$ form a basis of $\{\gamma^{\prime}(t)\}^{\perp}\subset T_{\mathcal{F}_{t}(x)}M$ along the entire geodesic $\gamma$.
We denote $A_{t}(x)$ to be the differential map:
$A_{t}(x):=D\mathcal{F}_{t}(x):T_{x}M\rightarrow T_{\mathcal{F}_{t}(x)}M.$
It is known that $A_{t}$ satisfies the Riccati equation
$D_{t}D_{t}A_{t}(x)+R(A_{t}(x),\gamma^{\prime}(t))\gamma^{\prime}(t)=0$
with the initial condition $A_{0}=\mathrm{Id}$ and $D_{t}A_{t}|_{t=0}=-\mathrm{Hess}\psi$, see page 142 of [20].
Since $\{J_{i}\}_{i\geq 2}$ form a basis of the vector space $\{\gamma^{\prime}(t)\}^{\perp}$, and since for $i\geq 2$, $\langle J_{i}^{\prime}(t),\gamma^{\prime}\rangle=0$ by Gauß lemma, we can define, as in [20], the matrix $U(t)$ to be such that $U(t)A_{t}:=D_{t}A_{t}$. And the Riccati equation can also be written as
$U^{\prime}+U^{2}+RA^{-1}=0\text{, or }U^{\prime}+U^{2}+R(\cdot,\partial_{t}\mathcal{F}_{t})\partial_{t}\mathcal{F}_{t}=0.$
At $\gamma(t)$ we set an orthonormal basis $\{e^{i}_{t}\}$ with $e^{1}_{t}:=\gamma^{\prime}(t)/\|\gamma^{\prime}(t)\|$ being the tangent vector of $\gamma(t)$ of unit length. In terms of this basis, we set $U_{ij}(t)=g(e^{i}_{t},U(t)e^{j}_{t})$, $\lambda_{t}=1+\int_{0}^{t}U_{11}(s)ds$, and $f(t)=\exp(\lambda_{t})$, the Riccati equation implies that
$\frac{d}{dt}U_{11}(t)+\sum_{j=1}^{n}U_{1j}^{2}=0\text{, therefore }-\lambda^{\prime\prime}=-\frac{d}{dt}U_{11}(t)=\sum_{j=1}^{n}U_{1j}^{2}\geq U_{11}^{2}=(\lambda^{\prime})^{2}$
@@
We let $y_{t}:=\mathrm{log}J_{t}$ where $J_{t}:=\mathrm{det}A_{t}=\|D\mathcal{F}_{t}(x)\|$, $y^{\psi}_{t}:=\mathrm{log}J^{\psi}_{t}=\psi(x)-\psi\circ\mathcal{F}_{t}(x)+y_{t}$ and $\alpha_{t}:=y^{\psi}_{t}-\lambda_{t}$. Define
@@
since $y^{\prime}=\mathrm{tr}U$, $y^{\prime\prime}=\mathrm{tr}(U^{\prime})$, we have
$\alpha^{\prime\prime}=y^{\prime\prime}-\lambda_{t}^{\prime\prime}-(\psi\circ\mathcal{F}_{t})^{\prime\prime}$ $=$ $-\mathrm{Ric}(\gamma^{\prime})-\mathrm{tr}(U^{2})-U_{11}^{\prime}-\mathrm{Hess}_{\psi}(\gamma^{\prime},\gamma^{\prime})$
$=$ $-\mathrm{Ric}(\gamma^{\prime})-\sum_{i,j=1}^{n}U_{ij}^{2}+\sum_{j=1}^{n}U_{1j}^{2}-\mathrm{Hess}_{\psi}(\gamma^{\prime},\gamma^{\prime})$
$\leq$ $-\mathrm{Ric}(\gamma^{\prime})-\sum_{i,j=2}^{n}U_{ij}^{2}-\mathrm{Hess}_{\psi}(\gamma^{\prime},\gamma^{\prime})$
$\leq$ $-\mathrm{Ric}(\gamma^{\prime})-\frac{1}{n-1}\Big(\sum_{i=2}^{n}U_{ii}\Big)^{2}-\mathrm{Hess}_{\psi}(\gamma^{\prime},\gamma^{\prime}).$
@@
Inequality (6) and the hypothesis $\mathrm{Ric}_{m,N}\geq K$ imply that $u$ satisfies the differential inequality, as in pages 143-144 of [20],
@@
Along [1], we can show a version of Borell-Brascamp-Lieb inequality on weighted Riemannian manifolds:
@@
The proof of the above theorem is essentially the same as the proof of the main theorem of [1], which is similar to the proof of the main theorem of [6] by replacing the volume distortion coefficient by the function $\beta^{t}_{K,N}$.
@@
The above theorem can be expressed in terms of the $p$-mean via the same method of scaling as the main theorem of [6] was generalised to Corollary 1.1 of [6].
@@ 3 Rigidity of the Borell-Brascamp-Lieb inequality
    The measure behaves as an (Nn)(N-n)-th power of a linear combination of trignometric functions along the geodesic $\gamma(t)$ when $K>0$; as an (Nn)(N-n)-degree polynomial of $t$ along $\gamma(t)$ when $K=0$; and as an (Nn)(N-n)-th power of a linear combination of hyperbolic functions along $\gamma(t)$ when $K<0$, for some constants $c>0$.
@@
We assume $N>n$ in the proof, as the case of $N=n$ reduces to [2]. With the above lemma and the construction of the functions $\alpha,y,\lambda_{t},u,f$ and the definition of the matrix $U$ as in the proof of Lemma 1, we have for all $t\in(0,1)$,
@@
The equality being attained has two implications. The first implication is that the matrix $V$, defined to be $V:=(U_{ij})_{i,j=2,...,n}$, satisfies the condition $\mathrm{tr}(V^{2})=\frac{1}{n-1}(\mathrm{tr}V)^{2}$, that it equals the scalar multiple of the identity matrix $V=\xi(t)\mathrm{Id}_{n-1}$, which further implies that the sectional curvatures along the geodesic $\gamma(t)$ are all identical by the Riccati equation.
@@
Thus, having $y^{\prime}-\lambda^{\prime}=\mathrm{tr}V=(n-1)\xi(t)$, we have $\partial_{\gamma^{\prime}}\psi=(N-n)\xi(t)$, and $\alpha^{\prime}=y^{\prime}-\lambda^{\prime}-\partial_{\gamma^{\prime}}\psi=(N-1)\xi(t)$. Hence equation (9) gives, when $\mathrm{Ric}_{m,N}=Kg$,
@@
$\frac{d^{2}}{dt^{2}}\psi\circ\gamma(t)-\frac{1}{N-n}\Big[\frac{d}{dt}(\psi\circ\gamma(t))\Big]^{2}=\frac{N-n}{N-1}Kr^{2}.$
Therefore, it can be deduced that $e^{-\psi\circ\gamma(t)}$ satisfies the differential equation
$\frac{d^{2}}{dt^{2}}e^{\frac{-\psi\circ\gamma(t))}{N-n}}=-\frac{Kr^{2}}{N-1}e^{\frac{-\psi\circ\gamma(t)}{N-n}}.$
@@
$e^{-\psi\circ\mathcal{F}_{t}(x)}=\left\{\begin{array}[]{cl}\Big[C_{0}\sin\Big(\sqrt{\frac{K}{N-1}}rt\Big)+C_{1}\cos\Big(\sqrt{\frac{K}{N-1}}rt\Big)\Big]^{N-n}\quad&\mbox{when $K>0$,}\\ (C_{0}t+C_{1})^{N-n}\quad&\mbox{when $K=0$,}\\ \Big[C_{0}\sinh\Big(\sqrt{\frac{-K}{N-1}}rt\Big)+C_{1}\cosh\Big(\sqrt{\frac{-K}{N-1}}rt\Big)\Big]^{N-n}\quad&\mbox{when $K<0$}\end{array}\right$.
where $C_{0}$ and $C_{1}$ are constants determined by the initial values $e^{-\psi(x)}$ and $\partial_{\gamma^{\prime}(0)}e^{-\psi(x)}$
$C_{0}=\frac{-\partial_{\gamma^{\prime}(0)}\psi(x)}{N-n}e^{\frac{-\psi(x)}{N-n}}\text{, }C_{1}=e^{\frac{-\psi(x)}{N-n}}.$
@@
For completeness, we include also a rigidity result on unweighted Riemannian manifolds, which is a direct consequence of Theorem 2 and can be considered as an addition of the consideration of dimension bounds to the curvature rigidity result in Theorem 4.1 of [2].

### 20 Apr 2026 v2

@@ Abstract
In this paper we discuss some results regarding the rigidity of the Borell-Brascamp-Lieb inequality and the Brunn-Minkowski inequality. We show a theorem of rigidity on curvature and measure of the Borell-Brascamp-Lieb inequality, a generalisation of the curvature rigidity theorem by Balogh-Kristály in (Balogh and Kristály, 2018) to the weighted setting. We present some rigidity results of the Brunn-Minkowski inequality and a few further open problems.
@@ 1 Introduction
Since the seminal works on the application of optimal transport in geometric analysis including (Otto and Villani, 2000), (Sturm, 2006) and (Lott and Villani, 2009), it became possible to study a wide range of analytical problems on Riemannian manifolds using the tool of optimal transport. In particular, the theory of optimal transport has been a powerful tool in the study of functional and geometrical inequalities on Riemannian manifolds with curvature and dimension bounds.
First proved in 1952 by Henstock and MacBeth on $\mathbb{R}$ by Henstock and Macbeath (1953) and later developed by Borell (1975) and Brascamp and Lieb (1976), the Borell-Brascamp-Lieb inequality is an integral inequality concerned with the $L^{1}$-norm of two functions and that of the mean of them.
@@
The Borell-Brascamp-Lieb inequality was proved on Riemannian manifolds by Cordero-Erausquin, McCann and Schmuckenschläger in (Cordero-Erausquin et al., 2001). It was then generalised further to metric measure spaces by Bacher (2010). The Borell-Brascamp-Lieb inequality implies a number of functional and geometric inequalities including the Brunn-Minkowski inequality and the Prékopa-Leindler inequality, a diagram showing their relationship can be seen in (Gardner, 2002).
Rigidity of the Borell-Brascamp-Lieb inequality on $\mathbb{R}^{n}$ has been studied since 1970s, various papers including (Dubuc, 1977) since then have discussed this problem. Rigidity of it on (unweighted) Riemannian manifolds with curvature bounded from below was shown by Balogh-Kristály in (Balogh and Kristály, 2018).
This paper generalises the curvature rigidity result of the Borell-Brascamp-Lieb inequality on Riemannian manifolds by Balogh and Kristály, Theorem 4.1 of (Balogh and Kristály, 2018), to weighted Riemannian manifolds. The main result of this paper states that the rigidity of the Borell-Brascamp-Lieb inequality implies identical sectional curvatures and certain behaviour of the measure.
@@
On a weighted Riemannian manifold $(M,g,m=\mathrm{e}^{-\psi}\mathrm{vol}_{g})$, satisfying $\mathrm{Ric}_{m,N}\geq Kg$ where $n\leq N<\infty$ and $K\in\mathbb{R}$, equality of the Borell-Brascamp-Lieb inequality implies that the $N$-Ricci curvature along the geodesic of transport $\gamma(t):=\mathcal{F}_{t}(x)$ equals $K$ and the sectional curvatures along the geodesic $\gamma$ are identically $\frac{K}{N-1}$. The measure behaves as an (Nn)(N-n)-th power of a linear combination of trignometric functions along the geodesic $\gamma(t)$ when $K>0$; as an (Nn)(N-n)-degree polynomial of $t$ along $\gamma(t)$ when $K=0$; and as an (Nn)(N-n)-th power of of a linear combination of hyperbolic functions along $\gamma(t)$ when $K<0$.
@@ 2 Preliminaries
Throughout the paper, we consider a smooth weighted Riemannian manifold $(M,g,m)$ of dimension $n$, where $m=\mathrm{e}^{-\psi}\mathrm{vol}_{g}$, satisfying the condition $\mathrm{Ric}_{m,N}(\cdot)\geq Kg(\cdot,\cdot)$, $N\in[n,\infty)$ and $K\in\mathbb{R}$. Recall that the $N$-Ricci curvature $\mathrm{Ric}_{m,N}$ is defined to be:
@@
By the theorem of Brenier-McCann, given two measures on $(M,g,m)$, $\mu_{0}=\rho_{0}m=\rho_{0}\mathrm{e}^{-\psi}\mathrm{vol}_{g}$ and $\mu_{1}=\rho_{1}m=\rho_{1}\mathrm{e}^{-\psi}\mathrm{vol}_{g}$ in $\mathcal{P}^{2}(M)$ that are absolutely continuous with respect to $m$, there is a unique optimal transference plan from $\mu_{0}$ to $\mu_{1}$, given by $\mathcal{F}_{t}$, $\mu_{t}=(\mathcal{F}_{t})_{\#}\mu_{0}$, the push-forward of $\mu_{0}$ by $\mathcal{F}_{t}(x):=\exp_{x}(t\nabla f(x))$, and $f:M\rightarrow\mathbb{R}$ is a $\frac{d^{2}}{2}$-convex function. We refer to (Villani, 2003) for the fundamentals of optimal transport theory, see also (McCann, 2001) and (Figalli and Gigli, 2011) for the theory of optimal transport on Riemannian manifolds.
We consider the weighted Jacobian $J^{\psi}_{t}(x):=\mathrm{e}^{\psi(x)-\psi(\mathcal{F}_{t}(x))}\|(D\mathcal{F}_{t}(x))\|$ along the geodesic $\gamma(t):=\mathcal{F}_{t}(x)$. We have the following concavity condition on the weighted Jacobian:
@@
The interpolation inequality (3) on unweighted Riemannian manifolds was shown in the proof of Theorem 1.7 of (Sturm, 2006), and its Finsler version was shown in the proof of Theorem 1.2 of (Ohta, 2009). Here we present a proof on weighted Riemannian manifolds satisfying $\mathrm{Ric}_{m,N}\geq Kg$, along the computation shown in the proof of Theorem 1.7 of (Sturm, 2006), with an addition of the factor due to measure to the proof in (Sturm, 2006), for later use. See also pages 393-401 of (Villani, 2009).
@@
On a weighted Riemannian manifold $(M,g,m=e^{-\psi(x)}\mathrm{vol}_{g})$ satisfying $\mathrm{Ric}_{m,N}\geq Kg$, the weighted Jacobian $J^{\psi}_{t}(x):=e^{\psi(x)-\psi(\mathcal{F}_{t}(x))}\|(D\mathcal{F}_{t}(x))\|$ defined along the geodesic $\gamma(t):=\mathcal{F}_{t}(x)$, where $\mathcal{F}_{t}(x)=\exp_{x}(t\nabla f(x))$ described as above satisfies the concavity condition (3).
@@
When $\gamma(t)$ is nonconstant, at $x$, we consider the tangent space $T_{x}M=T_{\mathcal{F}_{0}(x)}M$ and an orthonormal frame $\{e_{0}^{i}\}_{i=1}^{n}$, where we denote $e^{1}_{0}$ to be the unit tangent of the geodesic: $\frac{\gamma^{\prime}(0)}{\|\gamma^{\prime}(0)\|}$. We can thus define n1n-1 Jacobi fields $\mathbf{J}_{i}(t)=\frac{\partial}{\partial s}\exp_{x}(t(\nabla f(x)+se^{i}_{0}))$ for $i\in\{2,...,n\}$ that are orthogonal to $\gamma^{\prime}(t)$, and the vectors $\{\mathbf{J}_{i}(t)\}$ form a basis of $\{\gamma^{\prime}(t)\}^{\perp}\subset T_{\mathcal{F}_{t}(x)}M$ along the entire geodesic $\gamma$.
We denote $\mathrm{A}_{t}(x)$ to be the differential map:
$\mathrm{A}_{t}(x):=\mathrm{D}\mathcal{F}_{t}(x):T_{x}M\rightarrow T_{\mathcal{F}_{t}(x)}M.$
It is known that $\mathrm{A}_{t}$ satisfies the Riccati equation
$\mathrm{D}_{t}\mathrm{D}_{t}\mathrm{A}_{t}(x)+R(\mathrm{A}_{t}(x),\gamma^{\prime}(t))\gamma^{\prime}(t)=0$
with the initial condition $A_{0}=\mathrm{Id}$ and $\mathrm{D}_{t}A_{t}|_{t=0}=-\mathrm{Hess}\psi$, see page 142 of (Sturm, 2006).
Since $\{J_{i}\}_{i\geq 2}$ form a basis of the vector space $\{\gamma^{\prime}(t)\}^{\perp}$, and since for $i\geq 2$, $\langle\mathbf{J}_{i}^{\prime}(t),\gamma^{\prime}\rangle=0$ by Gauß lemma, we can define, as in (Sturm, 2006), the matrix $\mathbf{U}(t)$ to be such that $\mathbf{U}(t)\mathrm{A}_{t}:=\mathrm{D}_{t}\mathrm{A}_{t}$. And the Riccati equation can also be written as
$\mathbf{U}^{\prime}+\mathbf{U}^{2}+R\mathrm{A}_{t}^{-1}=0\text{, or }\mathbf{U}^{\prime}+\mathbf{U}^{2}+R(\cdot,\partial_{t}\mathcal{F}_{t})\partial_{t}\mathcal{F}_{t}=0.$
At $\gamma(t)$ we set an orthonormal basis $\{e^{i}_{t}\}$ with $e^{1}_{t}:=\gamma^{\prime}(t)/\|\gamma^{\prime}(t)\|$ being the tangent vector of $\gamma(t)$ of unit length. In terms of this basis, we set $\mathbf{U}_{ij}(t)=g(e^{i}_{t},\mathbf{U}(t)e^{j}_{t})$, $\lambda_{t}=1+\int_{0}^{t}U_{11}(s)ds$, and $f(t)=\exp(\lambda_{t})$, the Riccati equation implies that
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{U}_{11}(t)+\sum_{j=1}^{n}\mathbf{U}_{1j}^{2}=0\text{, therefore }-\lambda^{\prime\prime}=-\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{U}_{11}(t)=\sum_{j=1}^{n}\mathbf{U}_{1j}^{2}\geq\mathbf{U}_{11}^{2}=(\lambda^{\prime})^{2}$
@@
We let $y_{t}:=\mathrm{log}J_{t}$ where $J_{t}:=\mathrm{det}\mathrm{A}_{t}=\|\mathrm{D}\mathcal{F}_{t}(x)\|$, $y^{\psi}_{t}:=\mathrm{log}J^{\psi}_{t}=\psi(x)-\psi\circ\mathcal{F}_{t}(x)+y_{t}$ and $\alpha_{t}:=y^{\psi}_{t}-\lambda_{t}$. Define
@@
since $y^{\prime}=\mathrm{tr}\mathbf{U}$, $y^{\prime\prime}=\mathrm{tr}(\mathbf{U}^{\prime})$, we have
$\alpha^{\prime\prime}=y^{\prime\prime}-\lambda_{t}^{\prime\prime}-(\psi\circ\mathcal{F}_{t})^{\prime\prime}$ $=$ $-\mathrm{Ric}(\gamma^{\prime})-\mathrm{tr}(\mathbf{U}^{2})-\mathbf{U}_{11}^{\prime}-\mathrm{Hess}_{\psi}(\gamma^{\prime},\gamma^{\prime})$
$=$ $-\mathrm{Ric}(\gamma^{\prime})-\sum_{i,j=1}^{n}\mathbf{U}_{ij}^{2}+\sum_{j=1}^{n}\mathbf{U}_{1j}^{2}-\mathrm{Hess}_{\psi}(\gamma^{\prime},\gamma^{\prime})$
$\leq$ $-\mathrm{Ric}(\gamma^{\prime})-\sum_{i,j=2}^{n}\mathbf{U}_{ij}^{2}-\mathrm{Hess}_{\psi}(\gamma^{\prime},\gamma^{\prime})$
$\leq$ $-\mathrm{Ric}(\gamma^{\prime})-\frac{1}{n-1}\Big(\sum_{i=2}^{n}\mathbf{U}_{ii}\Big)^{2}-\mathrm{Hess}_{\psi}(\gamma^{\prime},\gamma^{\prime}).$
@@
Inequality (6) and the hypothesis $\mathrm{Ric}_{m,N}\geq K$ imply that $u$ satisfies the differential inequality, as in pages 143-144 of (Sturm, 2006),
@@
Along (Bacher, 2010), we can show a version of Borell-Brascamp-Lieb inequality on weighted Riemannian manifolds:
@@
The proof of the above theorem is essentially the same as the proof of the main theorem of (Bacher, 2010), which is similar to the proof of the main theorem of (Cordero-Erausquin et al., 2001) by replacing the volume distortion coefficient by the function $\beta^{t}_{K,N}$.
@@
The above theorem can be expressed in terms of the $p$-mean via the same method of scaling as the main theorem of (Cordero-Erausquin et al., 2001) was generalised to Corollary 1.1 of (Cordero-Erausquin et al., 2001).
@@ 3 Rigidity of the Borell-Brascamp-Lieb inequality
    The measure behaves as an (Nn)(N-n)-th power of a linear combination of trignometric functions along the geodesic $\gamma(t)$ when $K>0$; as an (Nn)(N-n)-degree polynomial of $t$ along $\gamma(t)$ when $K=0$; and as an (Nn)(N-n)-th power of a linear combination of hyperbolic functions along $\gamma(t)$ when $K<0$.
@@
We assume $N>n$ in the proof, as the case of $N=n$ reduces to (Balogh and Kristály, 2018). With the above lemma and the construction of the functions $\alpha,y,\lambda_{t},u,f$ and the definition of the matrix $U$ as in the proof of Lemma 1, we have for all $t\in(0,1)$,
@@
The equality being attained has two implications. The first implication is that the matrix $V$, defined to be $\mathbf{V}:=(\mathbf{U}_{ij})_{i,j=2,...,n}$, satisfies the condition $\mathrm{tr}(\mathbf{V}^{2})=\frac{1}{n-1}(\mathrm{tr}\mathbf{V})^{2}$, that it equals the scalar multiple of the identity matrix $\mathbf{V}=\xi(t)\mathbf{\mathrm{Id}}_{n-1}$, which further implies that the sectional curvatures along the geodesic $\gamma(t)$ are all identical by the Riccati equation.
@@
Thus, having $y^{\prime}-\lambda^{\prime}=\mathrm{tr}\mathbf{V}=(n-1)\xi(t)$, we have $\partial_{\gamma^{\prime}}\psi=(N-n)\xi(t)$, and $\alpha^{\prime}=y^{\prime}-\lambda^{\prime}-\partial_{\gamma^{\prime}}\psi=(N-1)\xi(t)$. Hence equation (9) gives, when $\mathrm{Ric}_{m,N}=Kg$,
@@
$\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\psi\circ\gamma(t)-\frac{1}{N-n}\Big[\frac{\mathrm{d}}{\mathrm{d}t}(\psi\circ\gamma(t))\Big]^{2}=\frac{N-n}{N-1}Kr^{2}.$
Therefore, it can be deduced that $\mathrm{e}^{-\psi\circ\gamma(t)}$ satisfies the differential equation
$\frac{\mathrm{d}^{2}}{\mathrm{d}t^{2}}\mathrm{e}^{\frac{-\psi\circ\gamma(t))}{N-n}}=-\frac{Kr^{2}}{N-1}\mathrm{e}^{\frac{-\psi\circ\gamma(t)}{N-n}}.$
@@
$\mathrm{e}^{-\psi\circ\mathcal{F}_{t}(x)}=\left\{\begin{array}[]{cl}\Big[C_{0}\sin\Big(\sqrt{\frac{K}{N-1}}rt\Big)+C_{1}\cos\Big(\sqrt{\frac{K}{N-1}}rt\Big)\Big]^{N-n}\quad&\mbox{when $K>0$,}\\ (C_{0}t+C_{1})^{N-n}\quad&\mbox{when $K=0$,}\\ \Big[C_{0}\sinh\Big(\sqrt{\frac{-K}{N-1}}rt\Big)+C_{1}\cosh\Big(\sqrt{\frac{-K}{N-1}}rt\Big)\Big]^{N-n}\quad&\mbox{when $K<0$}\end{array}\right$.
where $C_{0}$ and $C_{1}$ are constants determined by the initial values $\mathrm{e}^{-\psi(x)}$ and $\partial_{\gamma^{\prime}(0)}\mathrm{e}^{-\psi(x)}$
$C_{0}=\frac{-\partial_{\gamma^{\prime}(0)}\psi(x)}{N-n}\mathrm{e}^{\frac{-\psi(x)}{N-n}}\text{, }C_{1}=\mathrm{e}^{\frac{-\psi(x)}{N-n}}.$
@@
For completeness, we include also a rigidity result on unweighted Riemannian manifolds, which is a direct consequence of Theorem 2 and can be considered as an addition of the consideration of dimension bounds to the curvature rigidity result in Theorem 4.1 of (Balogh and Kristály, 2018).